OpenStudy (anonymous):

Calculus 1 limits!! Can someone check my work??? 1) lim (5x-3sinx0/(x) x-->0 The answer options are: a) 0 b) 8 c) 2 d) does not exist I got D but idk if that is right or if I did the right steps. Is that the same for anyone else? 2)lim (6x/2sinx) x-->0 My teacher did an example similar to this and I did this: [sin(9x)/2x] = 9/2 Is that correct? It's only the second day of summer classes so I'm trying to get the hang of these homework assignments as they are a bit different from his examples in class!

OpenStudy (anonymous):

\[\begin{align*}\lim_{x\to0}\frac{5x-3\sin x}{x}&=\lim_{x\to0}\frac{5x}{x}-\lim_{x\to0}\frac{3\sin x}{x}\\ &=5\lim_{x\to0}\frac{x}{x}-3\lim_{x\to0}\frac{\sin x}{x}\\ &=5\lim_{x\to0}1-3\lim_{x\to0}\frac{\sin x}{x}\\ &=5(1)-3(1)\\ &=2 \end{align*}\] \[\begin{align*}\lim_{x\to0}\frac{6x}{2\sin x}&=3\lim_{x\to0}\frac{x}{\sin x}\\ &=3(1)\\ &=3 \end{align*}\] Bonus: \[\begin{align*}\lim_{x\to0}\frac{\sin9x}{2x}&=\frac{9}{2}\lim_{x\to0}\frac{\sin9x}{x}\cdot\frac{1}{9}\\ &=\frac{9}{2}\lim_{x\to0}\frac{\sin9x}{9x}\\ &=\frac{9}{2} \end{align*}\] The takeaway here is to remember that \[\lim_{x\to0}\frac{\sin ax}{ax}=1\] for \(a\not=0\). The strategy is to rewrite the given expression into a form that resembles this and to keep track of your coefficients and added terms.